OpenStudy (anonymous):

In 5 years, Dad will be three times as old as his daughter Jill will be then. If the sum of their present ages is 50, how old are they now? If x + 5 is Jill's age five years from now, which of the following equations could be used to solve the problem? 3x + 5 = 55 - x 3x + 15 = 55 - x 3x + 15 = 45 - x

OpenStudy (anonymous):

Let x be the daughter's age and y be the dad's: $$ y + x = 50 \implies y = 50 - x \\ y + 5 = 3(x+5) \implies (50-x) + 5 = 3(x+5)\\ 55 - x = 3x + 15 \implies 3x + 15 = 55 - x $$

OpenStudy (anonymous):

\[ 3x + 15 = 55 - x \implies 4x = 40 \implies x = 10 \\ y = 50 - x \implies y = 50 - 10 = 40 \]